OpenStudy (anonymous):

What mass of ammonia gas is produced when 30.0 g of N2(g) and 10.0 g of H2(g) react according to the following (unbalanced) equation, N2(g) + H2(g) NH3(g)?

OpenStudy (julian.pitalua):

Balanced equation: N2(g) +3 H2(g) --->2NH3(g) moles(N2)= 30,0g/28,00g/mol=1,07 mol moles (H2)= 10,0g/2,00g/mol=5,00 mol calculating stoichiometric coefficients N2= 1,07/1= 1,07 H2= 5,00/3= 1,6 mol than we know that N2 is the limitant reactive Now we have to calculate number of moles(NH3) produced 1 mol (N2)/ 2 mol(NH3)= 1,07 mol (N2)/ xxx mol (NH3) I guess you already know the rest.